(y+x)p+(y-x)q=z


1
Expert's answer
2021-01-19T01:44:42-0500

The given partial differential equation is

(y+ x)p + (y− x)q = z

The Lagrange’s auxiliary equations are

"\frac{dx}{y+x}=\frac{dy}{y-x}=dz\\
\text{integrating the first two ratio, we have}\\
\frac{dx}{y+x}=\frac{dy}{y-x}\\
\frac{dy}{dx}=\frac{y-x}{y+x}\\
\text{solving by homogeneous first order differential equation}\\
y=vx\\
\frac{dy}{dx}=v+x\frac{dv}{dx}\\
v+x\frac{dv}{dx}=\frac{vx-x}{vx+x}\\
x\frac{dv}{dx}=\frac{v-1}{v+1}-v\\
x\frac{dv}{dx}=\frac{-(v^2+1)}{v+1}\\
\frac{dx}{x}=-\frac{v+1}{v^2+1}dv\\
\frac{dx}{x}+\frac{v+1}{v^2+1}dv\\=o\\

\int \frac{dx}{x}+\int \frac{v+1}{v^2+1}dv\\=c_1\\
c_1=log_ex+\frac{1}{2}log_e(v^2+1)+\frac{1}{v}tan^{-1}v\\
C_1=\frac{1}{2}log_e(\frac{x^3}{y}+1)+\frac{y}{x}tan^{-1}\frac{x}{y}\\
\text{integrating the last two ratio}\\
\frac{dy}{y-x}=\frac{dz}{z}\\
y=vz\\
\frac{dy}{dz}=v+z\frac{dv}{dz}\\
v+z\frac{dv}{dz}=\frac{vz-x}{z}\\
z\frac{dv}{dz}=\frac{vz-x}{z}-v\\
z\frac{dv}{dz}=\frac{-x}{z}\\
{z^2}dv=-xdz\\
\int {z^2}dv=\int -xdz\\
z^2v+xz=c_2\\
\frac{z^2y}{x}+xz=c_2\\
\phi(\frac{1}{2}log_e(\frac{x^3}{y}+1)+\frac{y}{x}tan^{-1}\frac{x}{y}, \frac{z^2y}{x}+xz)=0\\
\phi \\
\text{is an arbitrary constant}"